# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision Tasks

### v2

@@ 3.3 Dataset Analysis
Given the target images and their corresponding ICMscores, we now analyze how visual features of target images influence ICMscores. To investigate this relationship, we use a ResNet (He et al., 2015) model pre-trained on ImageNet (Deng et al., 2009) for object recognition tasks to extract feature representations for each image. We define the mean of all image features belonging to the same category as the centroid. Next, we compute the Pearson correlation (Sedgwick, 2012) between the L2 distances from its centroid to individual target image features and their ICMscores.
Figure 3: Correlation analysis between the feature distances of individual images to their categorical centroids and their corresponding ICMscore for the car and flower categories. Each “Car” and “Flower” image in ICMD is represented as a scatter point, with its corresponding ICMscore shown on the y-axis and the L2 distance from their centroids on the x-axis. The dashed lines, derived from linear regression, indicate a positive correlation between feature distance and CIMscore. As the L2 distance from the centroid increases, the ICMscore generally increases. The Pearson correlation coefficients are $r=0.46$ for “Car” and $r=0.42$ for “Flower” ($p<0.05$). See Sec. A.3 for the correlation analysis on other object categories.
Our results reveal a positive correlation between an image’s distance from its categorical centroid and its ICMscore across all object classes, as shown in Fig. 3 for the car and flower classes and Fig. A3 for the other object classes. The further away the features of an image are from its categorical centroid, the more memorable the image becomes. For example, a visually striking flower with distinct colors or intricate petal patterns stands out from typical flowers, making it more likely to be remembered (Fig. 1). Here, we explored the connections between memorability and feature distance within categories. Insights gained from this intra-class feature analysis motivate us to design the experiments that investigate the role of memorability in the following computer vision tasks in Sec. 4.
@@ 4.1 Memorability Prediction
Memorability prediction can provide insights into human cognition and perception by helping us understand how and why certain images are more memorable than others (Perera et al., 2019; Harada & Sakai, 2023; Needell & Bainbridge, 2022). Existing memorability prediction methods fail to disentangle the intrinsic and extrinsic factors that influence the memorability of an image. In contrast, our work focuses on predicting ICMscore based exclusively on intrinsic factors of images within the same category.
@@
Specifically, we use a ResNet (He et al., 2015), pre-trained on ImageNet (Deng et al., 2009) for image recognition tasks, to extract feature representations for an input image and then attach two fully connected layers to predict the memorability score. We train our ICMscore predictor on the ICMD dataset with an L2 regression loss, where the ICMscore of the input image serves as the ground truth. To validate the effectiveness of our predictor, we compare it against two baseline models. Baseline 1 employs the same network structure as our model but is trained from scratch without the pre-trained feature extractor on ImageNet for image recognition tasks. For Baseline 2, inspired by previous work (Hagen & Espeseth, 2023; Constantin & Ionescu, 2021), we use a Vision Transformer (ViT) (Alexey, 2020) trained from scratch. We perform 5-fold cross-validation to ensure the statistical robustness of the results.
Following the practices in memorability prediction (Isola et al., 2014; Hagen & Espeseth, 2023; Khosla et al., 2015), we assess the models’ performance by measuring the Spearman rank correlation (Zar, 2005) between the predicted and ground truth ICMscores. In Tab. 1, our model achieves the highest correlation (0.62 ± 0.06), demonstrating that ICMscores can be effectively predicted using features extracted from pre-trained image recognition models.
@@
To ensure the 0.62 correlation from our predictor is not due to model overfitting, we split human participants into two independent groups and computed the Spearman correlation of their ICMscores on the same target images. Across 10 runs, the average correlation was $0.69\pm 0.05$, demonstrating high consistency in within-human memorability. While our predictor’s correlation is slightly lower than human consistency, it surpasses both baselines, suggesting that features crucial for object recognition also contribute to predicting memorability. This finding aligns with the observed positive correlation between feature distance from the categorical centroids and ICMscores (Sec 3.3).
Table 1: Performance comparison of ICMscore prediction among different AI models. The Spearman Rank Correlation (Zar, 2005) quantifies the alignment between predicted and ground truth ICMscores. The higher, the better. The number of trainable parameters, the total number of model parameters, and the number of FLOPs (floating-point-operations) are also presented. Numbers in the brackets are standard errors.
Models Ours Baseline 1 Baseline 2
Spearman Rank Correlation 0.62(0.01) 0.59(0.2) 0.55(0.02)
Trainable Parameters 0.39M 21.68M 85.62M
Model Parameters 21.68M 21.68M 85.62M
FLOPs 3.68G 3.68G 16.86G
We detailed the number of trainable and total model parameters for our predictor and the two baselines in Tab. 1. Despite achieving accurate memorability predictions, our predictor has the fewest trainable parameters and the smallest model size, suggesting that larger models do not always yield better performance. Furthermore, our predictor requires the least number of FLOPs, demonstrating its high computational efficiency.
@@
To further explore what part of the image contributes to high memorability, we use Grad-CAM (Selvaraju et al., 2017) to visualize the importance maps for memorability, shown in Fig. 4. For HM images, our predictor focuses on distinctive regions. For instance, in a teapot image with an ICMscore of 0.78, our predictor emphasizes the unusual face texture of the teapot, while in a flower image with an ICMscore of 0.75, the model highlights the vibrant yellow flower in contrast to the smaller pale white flowers in the background. This suggests that exaggerated deviations from typical visual features make an image more memorable.
Figure 4: Visualization of the importance maps on two example images with high memorability highlighted by our ICMscore predictor based on Grad-CAM (Selvaraju et al., 2017). Grad-CAM visualizes key image regions by computing gradients of our ICMscore predictor against ground truth scores. For HM images, it highlights features that enhance memorability, with activation values shown in the color bar on the right.
@@ 4.2 Image Recognition
We employed ResNet (He et al., 2015) as the backbone with random weights, and trained it from scratch on each of the three curated subsets above using the Adam optimizer (Kingma & Ba, 2017). The initial learning rate was set to $1\times 10^{-3}$ and reduced by 50% every two epochs over a total of 10 epochs following a linear decay schedule. The resulting models were labeled as HM-model, MM-model, and LM-model based on their training subsets.
@@
We used the Cross-Entropy loss to train all three models for the 5-way classification task. Test images for object recognition were sourced from the original ImageNet (Deng et al., 2009) test set. However, as these test images lack ground truth ICMscores, we could not evaluate the top-1 accuracy of the models on subsets of test images stratified by ICMscores. Fig. 5 presents the average top-1 accuracies of the three models on all the test images for all 5 object classes over 5 random seeds.
@@
Our results in Fig. 5 indicate that the LM-model consistently outperforms the HM-model, achieving an average accuracy of $0.83$ compared to $0.77$ across all 5 object classes. Statistical analysis using one-way ANOVA revealed significant differences in top-1 accuracy between the two models ($p<0.05$). Surprisingly, the LM-model also surpasses the MM-model. These findings indicate that image memorability impacts image recognition, with LM images improving accuracy and HM images impairing it. We also provide the confusion matrices for the classification performances of all three models in Sec. A.4.
Figure 5: Performance comparison in Top-1 Accuracy between image classification models trained on HM, MM, and LM images. From the top to the bottom, Top-1 Accuracy is reported for models trained on LM images (blue), HM images (green) and MM images (yellow). The chance is 20% for the 5-way classification (red dash line). The error bar is the standard error calculated over 5 independent runs. One-way ANOVA across the results of all models yields p $<0.05$, indicating statistical significance in accuracy differences among green, yellow, and blue bars.
@@
In real-world applications, evaluating the memorability of all images through human behavioral studies is often impractical before training recognition models with images of varying memorability. Therefore, a computational model for predicting an image’s memorability becomes invaluable. We leveraged the ICMscore predictor (Sec. 4.1) to estimate the ICMscore for all images and then used these predictions as criteria to conduct the same experiment for image recognition. From Fig. A6 (a), consistent with findings above, recognition models trained exclusively on LM images categorized by predicted ICMscores achieved higher top-1 accuracies than the MM-model from our ICMD dataset. One possible reason is that easy images for object recognition often feature prototypical representations of object classes (Singh et al., 2023), which accelerate training and foster robust, generalizable learning. As shown in Sec 3.3, easy images tend to be nearer to categorical centroids, and hence, less memorable but more beneficial for learning.
@@
For comparison, we also conducted the same image recognition experiment using memorability definitions from (Bylinskii et al., 2015; Khosla et al., 2015; Dubey et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023) to divide the dataset into two subsets and utilizing the predictor from (Khosla et al., 2015). The results, shown in Fig. A6 (b), indicate that the predictor proposed by (Khosla et al., 2015) does not enhance image recognition accuracy. Additionally, we observed no significant difference in top-1 accuracy between models trained on the HM and LM images as predicted by (Khosla et al., 2015).
@@ 4.3 Continual Learning
Continual learning, particularly class-incremental learning, is essential for AI models to adapt to new classes over time without experiencing catastrophic forgetting of previously learned classes (Parisi et al., 2019). One type of continual learning methods is based on memory replay (Bagus & Gepperth, 2021; Tee & Zhang, 2023), which involves storing a subset of previously encountered examples from old classes in a memory buffer and replaying them during subsequent tasks. Given that some images are easier to remember while others are more easily forgotten, we conjecture that image memorability may play a key role in improving the continual learning performance of AI models. Here, we investigate how memorability can be leveraged to select exemplar images for replay, potentially enhancing the model’s ability to retain knowledge across tasks.
Figure 6: Continual learning performance of AI models using HM, MM, and LM images for replays. From left to right, the performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays is presented in Acc (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for the introduction to the evaluation metrics and the continual learning models. The error bars are the standard errors calculated over 5 independent runs with different random seeds. A two-tailed t-test analysis is performed to compare the performance in all three evaluation metrics between HM-and LM-models over 5 runs and between MM- and LM-models over 5 runs (p-values $<$ 0.05).
We structure our ICMD dataset for continual learning by training object recognition models on the first 2 object classes in the initial task, followed by introducing 1 incremental class in each subsequent task. Since there are a total of 5 object classes, this results in 4 tasks. To examine the role of image memorability, we further divide the images from each class into HM, MM, and LM image subsets by following the same procedures in Sec. 4.2. During replay, we exclusively used images from one of these subsets, meanwhile keeping training images identical across all 4 tasks. We label the continual learning models with replays from HM, MM, and LM images as HM-models, MM-models, and LM-models respectively. We train all these 3 continual learning models in the initial task for 10 epochs, and incremental tasks for 5 epochs with an Adam optimizer using a learning rate of $1\times 10^{-3}$. The replay buffer size is set to 20 images.
To benchmark the continual learning performance of these three models, we include the following three standard evaluation metrics, namely Accuracy at the last task (Acc), Backward knoWledge Transfer (BWT), and Forgetfulness (F). We define $A_{t,i}$ as the accuracy of the model trained on all the previous tasks up to task $t$ and tested on the test set of task $i$. Next, we introduce these metrics in details.
Accuracy at the last task (Acc) (Shi et al., 2024) is defined as the accuracy of the model trained up to the last task $T$ and tested on all the previous tasks. The higher the Acc, the better the model performs in continual learning tasks.
Backward Transfer (BWT)(Lin et al., 2022; Díaz-Rodríguez et al., 2018; Shi et al., 2024) quantifies the model’s knowledge retention ability by measuring performance drop on earlier tasks after learning new ones:
@@
Forgetting (F) (Shi et al., 2024) captures the accuracy drop on the test set of the first task for the model trained at the final task. It is defined as $F=A_{1,1}-A_{T,1}$. The lower the F, the better the model performs in continual learning tasks.
Results in Fig. 6 show that models using LM images for replay consistently outperform those using HM and MM images in Acc and BWT ($p<0.05$). Additionally, models using LM images exhibit reduced forgetting, indicated by lower F, though this difference is not statistically significant. We also plotted Acc and F as functions of task numbers for all three models in Fig. 7. The results indicate that the model replayed with LM images consistently exhibits reduced forgetting and achieves higher accuracy across all tasks. Together with our feature analysis in Sec. 3.3, our continual learning results align well with the previous research (Rebuffi et al., 2017; Ke et al., 2024) that replaying samples where their features are closer to the categorical centroids yields better continual learning performance.
Figure 7: Accuracy at the last task (Acc.) and forgetting (F) as a function of task numbers for continual learning models using HM, MM, and LM images for replays. Acc in solid lines and Forgetting in dotted lines at a given task are presented as a function of task numbers for HM-models (green), MM-models (yellow), and LM-models (blue). The y-axis (Score) could be interpreted differently as Acc scores (solid lines) or F scores (dotted lines) depending on the line types in the legend.
Similar to the image recognition experiments in Sec. 4.2, we categorized images into HM, MM, and LM using both our ICMscore predictor and the predictor from (Khosla et al., 2015). From Fig. A6, there is no performance difference in replay methods using LM and HM images predicted by (Khosla et al., 2015). In contrast, our ICMscore predictor remarkably reveals significant differences, with continual learning models replayed with LM images outperforming those with HM images or models using replay images predicted by (Khosla et al., 2015).
@@ 4.4 Memorability-controlled Image Editing
Editing images with controlled memorability offers valuable applications across various research fields, such as designing advertisements for commercial products and improving students’ knowledge retention ability in education (Goetschalckx et al., 2019; Garcia-Retamero et al., 2012; Moran et al., 2019). Next, we present a proof-of-concept demonstration of using a subset of teapot images in our ICMD dataset to train a state-of-the-art image diffusion model capable of modifying the memorability of an original teapot image.
To adjust a teapot image’s memorability while preserving its semantic content, we trained a high-memorability teapot generator (HM generator) using a subset of teapot images with high memorability (HM images) calculated from human behavioral experiments. First, for each HM image, we used SAM (Kirillov et al., 2023) to segment and mask the teapot area with white pixels. These masked images were paired with their corresponding original images to train image diffusion models. Specifically, we fine-tuned the pre-trained InstructPix2Pix model (Ip2p) (Brooks et al., 2022) with LoRA (Hu et al., 2021). The HM generator uses “highly memorable teapot” as the text prompt and the masked teapot image as input for image editing. The masked area guides the HM generator to focus on the specified region. During fine-tuning, the generator produces highly memorable teapot images and compares them against the ground truth HM images to improve its performance.
We qualitatively evaluated the HM generator’s performance using LM teapot images from our ICMD. As illustrated in Fig. 8, the generator effectively enhances memorability by adding visually striking textures to the original teapots. For quantitative evaluation, we used our memorability prediction model (detailed in Section 4.1) to predict ICMscores for both the original and generated images. The results indicate that the generated teapots consistently achieved higher ICMscores than the originals, demonstrating the effectiveness of our HM generator for memorability-controlled image editing.
We also employed a similar training strategy to create a Low-Memorability teapot generator (LM generator), with the results shown in Fig. 8. Similar analyses as applied to the HM generator were conducted. The findings reveal that memorability increases with the addition of distinctive visual elements, while it decreases when the teapot’s unique textures are removed.
@@
Figure 8: Proof-of-conept demonstration of memorability-controlled image editing. (a) During training, we fine-tune the image diffusion model, Ip2p  (Brooks et al., 2022), using LoRA (Hu et al., 2021) with pairs of masked images and their corresponding ground truth images. (b) During inference, SAM (Kirillov et al., 2023) segments and masks the teapot area of the source image, and our fine-tuned model takes the masked image as input to edit teapot images with varying memorability. (c) The high-memorability image generator increases the ICMscore of low-memorability (LM) teapots by adding striking textures and enhancing brightness. Conversely, the low-memorability image generator reduces the ICMscore by blurring distinctive features, making the teapots appear more ordinary with plain textures. The numbers beneath the images represent ICMscores predicted by our memorability predictor as introduced in Sec. 4.1.
5 Conclusion and Future Work

### v3

@@ 3.3 Dataset Analysis
Given the target images and their corresponding ICMscores, we now analyze how visual features of target images influence ICMscores. To investigate this relationship, we use a ResNet [35] model pre-trained on ImageNet [31] for object recognition tasks to extract feature representations for each image. We define the mean of all image features belonging to the same category as the centroid. Next, we compute the Pearson correlation [36] between the L2 distances from each centroid to individual target image features and their ICMscores.
Our results reveal a positive correlation between an image’s distance from its categorical centroid and its ICMscore across all object classes, as shown in Fig. 3(a). The further away the features of an image are from its categorical centroid, the more memorable the image becomes. Here, we explore the connections between memorability and feature distance within categories. Insights gained from this intra-class feature analysis motivate us to design the experiments that investigate the role of memorability in the following computer vision tasks in Sec. 4.
@@ 4.1 Memorability Prediction
Memorability prediction can provide insights into human cognition and perception by helping us understand how and why certain images are more memorable than others [37, 38, 39]. Existing memorability prediction methods fail to disentangle the intrinsic and extrinsic factors that influence the memorability of an image. In contrast, our work focuses on predicting ICMscore based exclusively on intrinsic factors of images within the same category.
@@
Specifically, we use a ResNet [35], pre-trained on ImageNet [31] for image recognition tasks, to extract image features, followed by two fully connected layers to predict memorability scores. We train our ICMscore predictor using the ICMD with an L2 regression loss, where the ICMscore of the input image serves as the ground truth. To validate the effectiveness of our predictor, we compare it against two baseline models. Baseline 1 employs the same network structure as our model but is trained from scratch without loading the pre-trained network parameters. For Baseline 2, inspired by previous work [30, 40], we use a Vision Transformer (ViT) pretrained on ImageNet [41], append two randomly initialized fully connected layers, and fine-tune only these added layers while keeping the ViT backbone frozen. We perform 5-fold cross-validation to ensure the statistical robustness of the results. Following the practices in memorability prediction [10, 30, 5], we assess the models’ performance by measuring the Spearman rank correlation [42] between the predicted and ground truth ICMscores. The first model achieves the highest correlation (0.64 ± 0.09), demonstrating that ICMscores can be effectively predicted using features extracted from pre-trained image recognition models, while the other two models obtained lower correlations: 0.60±0.2 and 0.57±0.05.
@@
To ensure the 0.64 correlation from our predictor is not due to model overfitting, we split human participants into two independent groups and computed the Spearman correlation of their ICMscores on the same target images. Across 10 runs, the average correlation is $0.69\pm 0.05$, demonstrating high consistency in within-human memorability. Although our model’s correlation is slightly lower than human consistency, it exceeds both baselines, suggesting that the model effectively captures diagnostic visual features of memorability rather than overfitting the data. This finding also aligns with the observed positive correlation between feature distance from the categorical centroids and ICMscores (Sec 3.3).
@@
To further explore what image components contribute to high memorability, we employ Grad-CAM [43] to visualize importance maps for memorability, as shown in Fig. 3(b). For highly memorable (HM) images, our predictor focuses on distinctive regions. For instance, in a teapot image with an ICMscore of 0.78, our predictor emphasizes the unique face-like texture of the teapot, while the model trained on LaMem [5] focuses on the teapot’s spout and handle which are typical features reflecting the generic categorical appearance of teapots.
@@ 4.2 Image Recognition
Figure 4: Performance of AI models using HM, MM, and LM images for image recognition task (a) and continual learning with replay (b) and (c). (a) Performance comparison in Top-1 Accuracy between image classification models trained on HM (green), MM (yellow), and LM (blue) images is presented. The chance level is 10% for the 10-way classification (red dashed line). Error bars represent standard errors calculated over 5 independent runs. One-way ANOVA across all models yields $p<0.05$, indicating statistically significant accuracy differences among green, yellow, and blue bars. (b) Performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays, is reported as Accuracy (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for details on evaluation metrics and continual learning models. Error bars represent standard errors calculated over 5 independent runs. Two-tailed t-test analysis comparing HM vs. LM models and MM vs. LM models across all three metrics shows significant differences ($p<0.05$). (c) Accuracy at the last task (solid lines) and Forgetting (dotted lines) as a function of task numbers for continual learning models using HM (green), MM (yellow), and LM (blue) images for replays. The y-axis (Score) represents either Accuracy or Forgetting as indicated in the legend.
@@
We employ ResNet [35] as the backbone with random weights, and train it from scratch on each of the three curated subsets above over all ten classes using the Adam optimizer [44]. The initial learning rate is set to $1\times 10^{-3}$ and reduced by 50% every five epochs over a total of 20 epochs following a linear decay schedule. The resulting models are labeled as HM-model, MM-model, and LM-model based on their training subsets.
@@
We use the Cross-Entropy loss to train all three models for the 10-way classification task. Test images for object recognition are obtained from the original ImageNet [31] test set regardless of their ICMscores. As these test images lack ground truth ICMscores, we could not evaluate the top-1 accuracy of the models on individual subsets of test images stratified by ICMscores. Fig. 4(a) presents the average top-1 accuracies of the three models on all the test images for all 10 object classes over 5 random seeds.
@@
Our results in Fig. 4(a) indicate that the LM-model consistently outperforms the HM-model, achieving an average accuracy of $0.80$ compared to $0.76$ across all 10 object classes. Statistical analysis using one-way ANOVA reveal significant differences in top-1 accuracy between the two models ($p<0.05$). Surprisingly, the LM-model also surpasses the MM-model. These findings indicate that image memorability impacts image recognition, with LM images improving accuracy and HM images impairing it. We also provide the confusion matrices for the classification performances of all three models in Sec. A.3.
@@
In real-world applications, training image recognition models in computer vision benefits from using images with varying memorability. However, obtaining memorability scores for all images through human behavioral studies is often time-consuming, costly, and therefore impractical. Therefore, a computational model for predicting an image’s memorability becomes invaluable. We leverage the ICMscore predictor, introduced in Sec. 4.1, to estimate the ICMscore for all images and then use these predictions as criteria to conduct the same experiment for image recognition. From Fig. A5 (a), consistent with the findings above, recognition models trained exclusively on LM images categorized by predicted ICMscores achieve higher top-1 accuracies than the MM-model. One possible reason is that easy images for object recognition often feature prototypical representations of object classes [45], which accelerate training and foster robust, generalizable learning. As shown in Sec 3.3, easy images tend to be nearer to categorical centroids, and hence, less memorable but more beneficial for representation learning in AI models.
@@
For comparison, we also conduct the same image recognition experiment based on the classical memorability definitions from prior works [4, 5, 6, 7, 8] to divide the dataset into two subsets and utilize the predictor from [5]. The results, shown in Fig. A5 (b), indicate that the predictor proposed by [5] does not enhance image recognition accuracy. Additionally, we observe no significant difference in top-1 accuracy between models trained on the HM and LM images as predicted by [5].
@@ 4.3 Continual Learning
Continual learning, particularly class-incremental learning, is essential for AI models to adapt to new classes over time without experiencing catastrophic forgetting of previously learned classes [46]. One type of continual learning methods is based on memory replay [47, 48, 45], which involves storing a subset of previously encountered examples in a memory buffer and replaying them during subsequent tasks. Given that some images are easier to remember while others are not, we conjecture that image memorability may play a key role in improving the continual learning performance of AI models. Here, we investigate how memorability can be leveraged to select exemplar images for replay, potentially enhancing the model’s ability to retain knowledge across tasks.
We structure our ICMD for continual learning by training object recognition models on the first 5 object classes in the initial task, followed by introducing 1 incremental class in each subsequent task. Since there are a total of 10 object classes, this results in 6 tasks. To examine the role of image memorability, we further divide the images from each class into HM, MM, and LM image subsets by following the same procedures in Sec. 4.2. During replay, we exclusively use images from one of these subsets, meanwhile keeping training images identical for all three continual learning models across all 6 tasks. We label the continual learning models with replays from HM, MM, and LM images as HM-models, MM-models, and LM-models respectively. We train all these 3 continual learning models in the initial task for 10 epochs, and incremental tasks for 5 epochs with an Adam optimizer using a learning rate of $1\times 10^{-3}$. The replay buffer size is set to 20 samples.
To benchmark the continual learning performance of these three models, we include the following three standard evaluation metrics, namely Accuracy at the last task (Acc), Backward knowledge Transfer (BWT), and Forgetfulness (F). We define $A_{t,i}$ as the accuracy of the model trained on all the tasks up to $t$ and tested on the test set of task $i$. Next, we introduce these metrics in details.
Accuracy at the last task (Acc) [49] is defined as the accuracy of the model trained up to the last task $T$ and tested on all the previous tasks. The higher the Acc, the better the model performs in continual learning tasks.
Backward Transfer (BWT)[50, 51, 49] quantifies the model’s knowledge retention ability by measuring performance drop on earlier tasks after learning new ones:
@@
Forgetting (F) [49] captures the accuracy drop on the test set of the first task for the model trained at the final task. It is defined as $F=A_{1,1}-A_{T,1}$. The lower the F, the better the model performs in continual learning tasks.
Results in Fig. 4(b) show that models using LM images for replay consistently outperform those using HM and MM images in Acc, BWT and F. We also plot accuracy and forgetting as functions of task numbers for all three models in Fig. 4(c). The results indicate that the model replayed with LM images consistently exhibits reduced forgetting and achieves higher accuracy across all tasks. Together with our feature analysis in Sec. 3.3, our continual learning results align well with the previous research [52, 53] that replaying samples where their features are closer to the categorical centroids yields better continual learning performance.
Similar to the image recognition experiments in Sec. 4.2, we categorize images into HM, MM, and LM using both our ICMscore predictor and the predictor from [5]. From Fig. A5, there is no significant performance difference in replay methods using LM and HM images predicted by [5]. In contrast, our ICMscore predictor remarkably reveals significant differences, with continual learning models replayed with LM images outperforming those with HM images or models using replay images predicted by [5].
@@ 4.4 Memorability-controlled Image Editing
Figure 5: Proof-of-conept demonstration of memorability-controlled image editing. (a) During training, we fine-tune the image diffusion model, Ip2p  [12], using LoRA [13] with pairs of masked images and their corresponding ground truth images. (b) During inference, SAM [54] segments and masks the teapot area of the source image, and our fine-tuned model takes the masked image as input to edit teapot images with varying memorability. (c) The high-memorability image generator increases the ICMscore of low-memorability (LM) teapots by adding striking textures and enhancing brightness. Conversely, the LM image generator reduces the ICMscore by blurring distinctive features, making the teapots appear more ordinary with plain textures. The numbers beneath the images represent ICMscores predicted by our memorability predictor as introduced in Sec. 4.1.
Editing images with controlled memorability offers valuable applications across various research fields, such as designing advertisements for commercial products and improving students’ knowledge retention ability in education [24, 55, 56]. Next, we present a proof-of-concept demonstration of using a subset of teapot images in our ICMD to train a state-of-the-art image diffusion model capable of modifying the memorability of an original teapot image.
To adjust a teapot image’s memorability while preserving its semantic content, we train a high-memorability teapot generator (HM generator) using a subset of teapot images with high memorability (HM images) calculated from human behavior experiments (Fig. 5a). First, for each HM image, we use SAM [54] to segment and mask the teapot area with white pixels. Next, these masked images are paired with their corresponding original images to train image diffusion models. Specifically, we fine-tuned the pre-trained InstructPix2Pix model (Ip2p) [12] with LoRA [13]. The generated images by Ip2p are compared against the ground truth HM images to improve its image generation quality.
During inference, the HM generator uses “highly memorable teapot" as the text prompt and the masked teapot image as input for image editing (Fig. 5c). The masked area guides the HM generator to focus on the specified region. The generator produces highly memorable teapot images. We qualitatively evaluate the HM generator’s performance using LM teapot images from our ICMD. As illustrated in Fig. 5c, the generator effectively enhances memorability by adding visually striking textures to the original teapots. For quantitative evaluation, we use our memorability prediction model (detailed in Section 4.1) to predict ICMscores for both the original and generated images. The results indicate that the generated teapots consistently achieve higher ICMscores than the originals, demonstrating the effectiveness of our HM generator for memorability-controlled image editing.
We also employ a similar training strategy to create a Low-Memorability teapot generator (LM generator), with the results shown in Fig. 5. Similar analyses as applied to the HM generator are conducted. The findings reveal that memorability increases with the addition of distinctive visual elements, while it decreases when the teapot’s unique textures are removed.
5 Discussion